Grade 12th passGeneral Physics

an engineer works at a plant out-of-town. a car is sent for him from the plant everyday that arrives at the railway station at the same time as the train he takes. one day the engineer arrived at the station one hour before his usual time and , without waiting for the car ,started walking for work . on his way he met the car and reached the plant 10 minutes before the usual time. how long did the engineer walk before he met the car ?the car moves with same speed everyday.

Kamalesh
8 Years agoGrade 12th pass

1 Answer

Saurabh Koranglekar
6 Years ago
Usual day :
Car starts fromFFatt=0t=0, reaches station atTTand again reaches at the factory at time2T2T.
Person reaches 'S' atT−60T-60. Car starts att=0t=0fromFF. Person walks for timettand reaches pointPPat timeT−60+tT-60+t. At this time car also reaches 'P'. Car comes back at 'F' at time(2T−10)(2T-10). That means car takes timet−5t-5fromF→PF→P. That means car reach at 'P' at timeT−5T-5.
NowT−5=T−60+t⇒t=55min